5 digits
100 digit is 2 more then my tens digit, my ones digit is 4 more then my thousands, the sum of my number is 8

Answer by ankor@dixie-net.com(22740)

5 digits: a, b, c, d, e
:
100 digit is 2 more then my tens digit,
c = d+2
:
my ones digit is 4 more then my thousands,
e = b+4
:
the sum of my number is 18
a + b + c + d + e = 8
Substitute:
a + b + (d+2) + d + (b+4) = 18 (clearly a sum of 8 is impossible)
a + 2b + 2d + 6 = 18
a + 2b + 2d = 18 - 6
a + 2b + 2d = 12
How about a=2, b=3, d=2
then
a=2
b=3
c=4
d=2
e=7
the number: 23427 is one possibility